What is the cost of buying 6kg of nuts at €3.00 per 500g ?
Answer by stanbon(75887)

What is the cost of buying 6kg of nuts at €3.00 per 500g ?
--------
6kg = 6000 g
------
Cost = 3*6000/500 = 3*12 = $36.00
--------------
Cheers,
Stan H.
--------------